Suppose the restriction enzyme GsuII binds a site that is 2 nucleotides in length. How often (how many base pairs) does this enzyme cut DNA?

Your answer must be numeric.
Answer #1

Number of different types of bases in DNA = 4

The formula for calculating the probability for finding a restriction site = 1/4n
n = number of bases in the recognition sequence

If a restriction enzyme recognizes a dinucleotide sequence, n = 2
= 1/42
= 1/16

There is a theoretical probability of finding a recognition site for every 16 nucleotides.

Please provide a POSITIVE RATING. Thank you